No change in the status of the POES spacecraft.
NOAA-14 rev 32832 / F at 0300Z on
May 14 - GAC1B noisy from spacecraft with 98.7% of the expected data
Additional recovery attempts not possible due to recorder scheduling.
POES operations were nominal over the past 72 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 72 hours.
No change in the status of the GOES spacecraft.
At 2145Z, 2226Z and 2245Z on May 12, GOES-8
imager frame breaks were observed and attributed to local thunderstorms
At 2230Z on May 12 on GOES 10 the PM operator
reports mismatch of imager and sounder IMC ID (F000 & F001
respectively). Investigation reveals problem is isolated to the SPS. CDA notified and their investigation reveals possible corruption during IMC update period due to the SPS ingesting bad data during thunderstorms. SOCC investigation also detected landmarks are being rejected by OATS due to mismatch of PM and OATS IMC IDís. After manually generating the IMC ID and cycling SPS to idle and back again at CDA, image frame start for the PACUS image corrected IMC ID (first frame start after CDA cycled SPS to idle at 2341Z) and landmarks ingested by OATS started at 2345Z.
Between 2102Z and 2337Z on May 12, GOES-10
numerous imager frame breaks were observed and attributed to local
thunderstorms at CDA.
GOES operations were nominal over the past 72
Questions or comments to Tim.Walsh@noaa.gov